# BloombergGPT: A Large Language Model for Finance

### v1

@@ Abstract
The use of NLP in the realm of financial technology is broad and complex, with applications ranging from sentiment analysis and named entity recognition to question answering. Large Language Models (LLMs) have been shown to be effective on a variety of tasks; however, no LLM specialized for the financial domain has been reported in literature. In this work, we present BloombergGPT, a 50 billion parameter language model that is trained on a wide range of financial data. We construct a 363 billion token dataset based on Bloomberg’s extensive data sources, perhaps the largest domain-specific dataset yet, augmented with 345 billion tokens from general purpose datasets. We validate BloombergGPT on standard LLM benchmarks, open financial benchmarks, and a suite of internal benchmarks that most accurately reflect our intended usage. Our mixed dataset training leads to a model that outperforms existing models on financial tasks by significant margins without sacrificing performance on general LLM benchmarks. Additionally, we explain our modeling choices, training process, and evaluation methodology. As a next step, we plan to release training logs (Chronicles) detailing our experience in training BloombergGPT.
@@ 2 Dataset
2.1 Financial Datasets (363B tokens – 54.2% of training)
@@
The Colossal Clean Crawled Corpus (C4) is a common dataset used to train LLMs, and was introduced to support training T5 (Raffel et al., 2020). Although it overlaps with Pile-CC, C4 is cleaned and processed differently; hence, we feel that including C4 in addition to The Pile can add value more than duplicated documents would. We find that C4 contains high-quality natural language documents due to the layers of cleaning, though others have noted that the distribution across web domains is unusual, with a high fraction of data stemming from patents Dodge et al. (2021).
@@ 3.1 Architecture
Our model is a decoder-only causal language model based on BLOOM (Scao et al., 2022). We present an overview of the architecture, with full details in § A.
@@ 4 Training Run
The process of training BloombergGPT involved decisions along the way based on the progress of model training. We share some highlights of this process. Figure 2 shows the learning curves for both training and validation sets. The solid lines show (smoothed) training loss and the dotted lines show loss on the held-out validation set. Changes in the color of the lines indicate changes to the optimization hyperparameter configurations, either as scheduled, or in response to increasing or stagnating validation loss. This plot shows the path taken by the successful model training run. To present a clear plot, the Figure does not show other attempts with different model configurations, overwritten partial runs after a rollback, or other training strategies not utilized in the final model.
@@ 7 Related Work
Transformer-based models rely on positional embeddings to encode position and location information of words in a text. Encoding the sequence position and the effect of this choice on model performance have been studied extensively. These include sinusoidal embeddings (Vaswani et al., 2017), rotary position embeddings (Su et al., 2021), adding relative position bias (Raffel et al., 2020), and adding linear biases to attention heads (Press et al., 2022). A side-effect of the strategy in Press et al. (2022) is that one can train on shorter sequences without loss in performance on longer sequences. This has two benefits: first, models learn to generalize (extrapolate) to longer sequences and second, models can be trained on shorter sequences reducing training time.
@@ 8.2 Openness
Nevertheless, our insights and experiences in training and evaluating BloombergGPT contribute to the developing understanding of these models. In particular, our experience may be useful to those building their own domain-specific models. During the process of developing BloombergGPT, we found the OPT chronicles, experiences of the BLOOM team, as well as work of non-open models like GPT-3, PaLM, Chinchilla, Galactica, and Gopher, to be crucial enablers of our work.
@@ A.1 Full Architecture
$\mathbb{P}(x_{j+1}=w|\{x_{t}\}_{t=1}^{j})=\mathop{\mathrm{softmax}}\nolimits(\boldsymbol{y_{j}})_{w},$ (6)
@@ A.2 SelfAttention with ALiBi ($\mathop{\mathrm{SA}}\nolimits$)
Let $n\in\{1,\mathellipsis,N\}$ denote an attention head where $N$ is the total number of heads. Let $D^{n}$ denote the dimensionality of each head. Let $\boldsymbol{A^{n}},\boldsymbol{M}\in\mathbb{R}^{T\times T}$ denote the ALiBi matrix and the attention mask, respectively, which will be defined later.
@@ A.5 List of All Trainable Parameters
B Details on external financial tasks

### v2

@@ Abstract
The use of NLP in the realm of financial technology is broad and complex, with applications ranging from sentiment analysis and named entity recognition to question answering. Large Language Models (LLMs) have been shown to be effective on a variety of tasks; however, no LLM specialized for the financial domain has been reported in literature. In this work, we present BloombergGPT, a 50 billion parameter language model that is trained on a wide range of financial data. We construct a 363 billion token dataset based on Bloomberg’s extensive data sources, perhaps the largest domain-specific dataset yet, augmented with 345 billion tokens from general purpose datasets. We validate BloombergGPT on standard LLM benchmarks, open financial benchmarks, and a suite of internal benchmarks that most accurately reflect our intended usage. Our mixed dataset training leads to a model that outperforms existing models on financial tasks by significant margins without sacrificing performance on general LLM benchmarks. Additionally, we explain our modeling choices, training process, and evaluation methodology. We release Training Chronicles (Appendix C) detailing our experience in training BloombergGPT.
@@ 2 Dataset
2.1 Financial Datasets (363B tokens – 51.27% of training)
@@
The Colossal Clean Crawled Corpus (C4) is a common dataset used to train LLMs, and was introduced to support training T5 (Raffel et al., 2020). Although it overlaps with Pile-$C$, C4 is cleaned and processed differently; hence, we feel that including C4 in addition to The Pile can add value more than duplicated documents would. We find that C4 contains high-quality natural language documents due to the layers of cleaning, though others have noted that the distribution across web domains is unusual, with a high fraction of data stemming from patents Dodge et al. (2021).
@@ 3.1 Architecture
Our model is a decoder-only causal language model based on BLOOM (Scao et al., 2022). We present an overview of the architecture, with full details in Appendix A.
@@ 4 Training Run
The process of training BloombergGPT involved decisions along the way based on the progress of model training. We share some highlights of this process. A detailed presentation appears in the Training Chronicles (Appendix C). Figure 2 shows the learning curves for both training and validation sets. The solid lines show (smoothed) training loss and the dotted lines show loss on the held-out validation set. Changes in the color of the lines indicate changes to the optimization hyperparameter configurations, either as scheduled, or in response to increasing or stagnating validation loss. This plot shows the path taken by the successful model training run. To present a clear plot, the Figure does not show other attempts with different model configurations, overwritten partial runs after a rollback, or other training strategies not utilized in the final model.
@@ 7 Related Work
Transformer-based models rely on positional embeddings to encode position and location information of words in a text. Encoding the sequence position and the effect of this choice on model performance have been studied extensively. These include sinusoidal embeddings (Vaswani et al., 2017), rotary position embeddings (Su et al., 2021a), adding relative position bias (Raffel et al., 2020), and adding linear biases to attention heads (Press et al., 2022). A side-effect of the strategy in Press et al. (2022) is that one can train on shorter sequences without loss in performance on longer sequences. This has two benefits: first, models learn to generalize (extrapolate) to longer sequences and second, models can be trained on shorter sequences reducing training time.
@@ 8.2 Openness
Nevertheless, our insights and experiences in training and evaluating BloombergGPT contribute to the developing understanding of these models. In particular, our experience may be useful to those building domain-specific models. During the process of developing BloombergGPT, we found the OPT chronicles, experiences of the BLOOM team, as well as work of non-open models like GPT-3, PaLM, Chinchilla, and Gopher, to be crucial enablers of our work. In support of this tradition, we include our Training Chronicles (Appendix C).
@@ A.1 Full Architecture
$\mathbb{P}(x_{j+1}=w|\{x_{t}\}_{t=1}^{j})=\mathop{\mathrm{softmax}}\nolimits(\boldsymbol{y_{j}})_{w}$ (6)
@@ A.2 SelfAttention with ALiBi ($\mathop{\mathrm{SA}}\nolimits$)
Let $n\in\{1,\mathellipsis,N\}$ denote an attention head where NN is the total number of heads. Let $D^{n}$ denote the dimensionality of each head. Let $\boldsymbol{A^{n}},\boldsymbol{M}\in\mathbb{R}^{T\times T}$ denote the ALiBi matrix and the attention mask, respectively, which will be defined later.
@@ A.5 List of All Trainable Parameters
Appendix B Details on external financial tasks
@@
Appendix C Training Chronicles
C.0 Still [Uncaptioned image]
Figure 7: Learning curve of our first training attempt named v0. Observe the large gap between training and validation losses, as well as the flatness of both curves after step 20k. The final 6k steps lasted about $\raisebox{2.15277pt}{\texttildelow}2.3$ days.
Our first training run was called v0. In this run, we experimented with curriculum learning. Data that the model would see in the future would likely be similar to the newer data in our training corpus, so we wanted the model to do better on those future documents. Additionally, since there are facts that change over time, newer information should ideally override the old. Therefore, we temporally ordered the training data by month in FinPile.
Figure 7 shows the learning curve for run v0. We observed a large gap between training and validation losses, which was expected: early stages of training would observe the oldest data (starting from 2007) whereas our validation set was strictly from the future (i.e., 2022). However, one week into training we found the model stuck on both training and validation loss, as seen by the very limited validation progress between steps 15k-20k and almost no progress after step 20k. There was the possibility that the training loss and the divergence of training and validation loss would both resolve themselves as the training data became more and more similar to the validation data as the curriculum progressed. However, we deemed this to be too risky to catch any other potential problems with the training that might require early intervention, since it would mean training for many steps without any diagnostic signal. We thus decided to abandon curriculum learning altogether.
We removed curriculum learning by shuffling all of our training data uniformly on the shard level. [^3] Instead of loading one shard of data at a time, we load multiple random shards (without replacement) at the same time and shuffle them on the fly. We then started a new run (v1.0), which led to much faster improvements in the validation loss. We were unable to ascertain if curriculum learning had a negative impact on training or if the loss plateaued due to other factors, for example, the other discovered issue in v1.x.
Figure 8: Gradient norms (top) and train & validation loss (bottom) of v1.x runs.
C.1 Elbow [Uncaptioned image]
During our new run without curriculum learning (v1.0), we observed that the gradient norm showed a steady increase after about 12k steps (~4.5 days of training), with occasional spikes (see Figure 8). This was accompanied by sudden jumps in the validation loss, possibly indicating that the model might be becoming sensitive to small changes in its weights. Training loss seemed to have been plateauing again, as well.
We believed that the gradient norm increases were the cause of the validation loss problems (notice the alignment between sudden validation loss jumps with some of the sudden gradient norm jumps for v1.0, in Figure 8). We made several attempts across several model runs to fix the gradient norm increases:
Run Changes from v1.0 run
Shared Change - Fully shuffle any data not seen by the model checkpoint that we chose to start (or restart) from instead of shard-level shuffling
v1.1 - Start from step 10.8k of v1.0, prior to any gradient increase
- Reduce max learning rate from 1e-4 to 8e-5
v1.2 - Continue from step 20.1k of v1.1 (most recent checkpoint)
- Reduce max learning rate from 1e-4 to 6e-5
- Reduce gradient clip from 1.0 to 0.3
v1.3 - Start from step 10.8k of v1.0, prior to any gradient increase
- Use FP32 precision for LM-head computation (prior to softmax)
v1.4 - Start from step 12.0k of v1.3
- Reduce max learning rate from 1e-4 to 6e-5
- Reduce gradient clip from 1.0 to 0.3
- Use FP32 precision for LM-head computation (prior to softmax)
Figure 9: Rescaled norms for each component in v1.0 run. Input LayerNorm at Layer 1 stood out.
Figure 10: Values for Input LayerNorm at Layer 1 in v1.0 run.
All of these attempted fixes were made after we observed a trend of increasing gradient norms similar to the original run (v1.0), or some early signs of a similar path that we hypothesized would eventually grow more. Since we didn’t want to waste training time, we did our best to make decisions early instead of allowing the model to continue down a bad training path.
We investigated the norms of the weights themselves to see if any peculiar trends were aligning with the gradient growth. In particular, we were curious to see if there were particular layers or components that were responsible for the large gradient norms.
Figure 9 plots L2 norms for each component, averaged by the square root of the number of elements (layer norm multipliers start from 1 and all the others start close to zero due to initialization). We observed that all components follow a similar benign trend except one: Input LayerNorm at layer 1 (i.e. $\mathop{\mathrm{LN}}\nolimits^{in}_{1}$), which suddenly elbows and starts increasing roughly linearly after step ~12k. This also aligns with the initial growth of the gradient norms.
To take a closer look, we inspected individual values of the multiplier weights $\gamma^{in}_{1}$ (there are 60 such values in a single model shard out of 128) in Figure 10. We observed all values contributing to the same trend of shrinking until steps 11-12k and then shifting to move upward instead.
During this investigation, we discovered another bug: Weight decay was applied to all the non-bias parameters, as opposed to skipping the LayerNorm multiplier weight $\gamma$ since they are initialized at 1. To the best of our knowledge, this practice has been inherited from the BERT implementation [^4] https://github.com/google-research/bert/blob/eedf5716ce1268e56f0a50264a88cafad334ac61/optimization.py#L59-L65.
This makes the elbow artifact shown in Figure 10 even more confusing: An additional push of weights towards 0 would trivially explain a downward trend but not a sudden shift to growth.
After four failed attempts to fix the run, we considered the possibility of this run being unsalvageable and contemplated starting from scratch to apply a more conservative hyperparameter setting from the beginning. These would include things that we have tried in our attempts such as shrinking the learning rate or gradient clipping. Additionally, because the pathological trend change is isolated to $\mathop{\mathrm{LN}}\nolimits^{in}_{1}$, which is topologically very close to the removed LayerNorm at the embedding layer ($\mathop{\mathrm{LN}}\nolimits^{em}$) we decided to add back $\mathop{\mathrm{LN}}\nolimits^{em}$ as an additional precaution, despite most other LLMs not having this component.
C.2 Slide [Uncaptioned image]
Figure 11: Loss, gradient norm and weight norms for listed components for the v2.0 run.
After numerous attempts to fix the elbow issue, we wanted to be as conservative as possible for the hyperparameter choices when starting from scratch to keep the learning dynamics as stable as possible. We started the next run (v2.0) with the following hyperparameters and changes:
  • Use FP32 precision in LM-head ($\mathop{\mathrm{softmax}}\nolimits$ in Equation 6)
  • Use max learning rate of 6e-5 instead of 1e-4
  • Use a gradient clipping value of 0.3 instead of 1.0
  • Fully shuffle data
  • Use a different seed to ensure different initialization and data order
  • Reintroduce LayerNorm at embedding layer ($\mathop{\mathrm{LN}}\nolimits^{em}$)
  • Use a longer learning rate warm-up period of 1800 steps
[…]
  • Use Megatron initialization rescaling (see use of $z^{\prime}$ in § A.5)
  • Apply query_key_layer_scaling (Shoeybi et al., 2019)
  • Apply a batch size warm-up: Use a batch size of 1024 for 7200 iterations, then increase to 2048
In addition to hyperparameter changes, we also performed additional monitoring to catch issues earlier. Because we observed the pathological behavior at the first LayerNorm component, we started monitoring the norms of the weights $\gamma^{em}$ and $\gamma^{in}_{1}$ (scaled by $1/\sqrt{D}$).
With the aforementioned conservative choice of hyperparameters during v2.0, we observed very smooth and (thankfully!) uneventful training for approximately 42 days (~115,500 iterations). We saw few surprises both in terms of training and validation performance curves (see Figure 11), as well as the norms of the gradients. The only intervention needed during this period was to restart the job after 28 days, due to the underlying platform having a hard limit on the duration of the job.
During this period, we observed a smoothly decreasing validation loss (except a few jumps earlier on) until it started to flatten around 2.116 (y-axis) at the end. Running training loss has a similar trend of overall decrease with the typical jitter and random occasional increments.
We also observed that weight norms for LayerNorm components of the initial layer were smooth and stable without any immediate or long-term trend changes (see Figure 11). This presents some evidence that we were not suffering from the pathological behavior we observed in v1.x in regards to LayerNorm parameters $\gamma^{em}$.
Overall, while this set of changes led to a smooth training run in v2.0, we cannot conclude which of these changes was decisive in leading to a successful training run. We defer such investigation to future work.
C.3 Suspense [Uncaptioned image]
Figure 12: Loss values for various runs at the end of v2. Note v2.4 overlaps with v2.6.
About 48 days into training v2.0, we noticed that the validation loss had not improved in a week (from iteration 115,500 to 133,200, see Figure 12, v2.0 curves). During the same period, we also noticed training loss flattening around 2.10 (with the usual jitter). We suspected that the model was no longer learning properly, and decided to intervene.
We considered two options: 1) changing the max learning rate, 2) rolling back to an earlier checkpoint and re-shuffling the remainder of the data to pick up a different optimization path.
We had two proposed ways in which to change the learning rate. An argument for increasing the learning rate was the possibility that we were stuck in a local optimum. Allowing the optimizer to make bigger jumps would allow the model to escape the optimum and continue learning. On the other hand, the argument for decreasing the learning rate was based on Zhang et al. (2022a) in which they had observed improvements after shrinking the learning rate after getting stuck. Furthermore, we had spent more steps in the high-learning-rate region of the overall learning rate schedule since, by following the Chinchilla scaling law, we had more total steps compared to models like BLOOM or GPT-3.
The other option was to roll back to an earlier checkpoint, re-shuffle the remainder of the data and continue training. Chowdhery et al. (2022) found that when they saw spikes in the validation loss, they “re-started training from a checkpoint roughly 100 steps before the spike started, and skipped roughly 200–500 data batches.” This suggests that data ordering mattered, and backing out of a bad data/gradient path may help. That may have been our issue with curriculum learning (v0.x), although it may have been that the issues were not with curriculum learning but with other issues we fixed in v1.0.
In the end, we decided to shrink the learning rate, roll back to the start of the increasing validation loss trend  7 days prior, and also re-shuffle the remaining data.
[…]
v2 v2.1 v2 v2.1 v2 v2.1 v2 v2.1
[…]
99300 8.43 8.60 37.77 43.57
115500 8.30 8.43 38.79 43.10
126600 8.34 8.24 8.40 8.32 37.86 38.09 42.37 42.79
133200 8.35 8.38 37.02 42.26
[…]
Using our two dev sets and downstream evaluations for guidance, we made several attempts to improve run v2.0 and direct the model to continue learning. A summary of our attempts follows: Run Changes from v2.0 run Shared Change - Re-shuffle future data starting from step 115500 v2.1 - Start from v2.0 step 115500
- Reduce max learning rate from 6e-5 to 4e-5
v2.2 - Start from v2.0 step 115500
- Increase dropout from 0.0 to 0.1
v2.3 - Start from v2.1 step 129900
- Reduce max learning rate from 6e-5 to 2e-5
- Increase dropout from 0.0 to 0.1
v2.4 - Start from v2.1 step 129900
- Reduce max learning rate from 6e-5 to 2e-5
v2.5 - Start from v2.3 step 137100
- Reduce max learning rate from 6e-5 to 1e-5
- Increase dropout from 0.0 to 0.1
v2.6 - Start from v2.1 step 129900
- Reduce max learning rate from 6e-5 to 2e-5
- Reduce weight decay from 0.1 to 0.01
After we lowered the learning rate and rolled back the model (v2.1), we observed an initial sudden (and dramatic) improvement; however, validation loss quickly flattened out. Coupled with the mixed results on downstream evaluation, we decided to enable dropout for the first time with a probability of 0.1.
With dropout, as expected, we observed a larger training loss since dropout is applied during the computation of the loss (v2.2 in Figure 12). However, we observed an initially decreasing validation loss. Still, as the run progressed further, validation loss started creeping back up and met the value of the original run (v2.0, blue).
Based on these observations, we decided that further decreasing the learning rate would give us the best chance to continue learning successfully. We subsequently tried various combinations of smaller values of learning rate and adding dropout. Observe that in v2.3 (red) with 2e-5 max learning rate and in v2.5 (brown) with 1e-5 as its continuation, both with a dropout rate of 0.1, shown in Figure 12. In Table 20, we observed v2.3 led to much better perplexity and slightly better downstream performance, and v2.5 continues to improve downstream performance compared to v2.3 in the beginning, while decreasing perplexity slightly. v2.4 (purple) attempted a max learning rate of 2e-5 as well, without dropout, however. The only odd run during this time is v2.6 (pink), in which we experimented with a smaller weight decay of 0.01 (compared to the original 0.1) with a max learning rate of 2e-5 to investigate the possibility of getting stuck in local minima due to too strong of a pull from the weight decay. However, this yields almost the exact same curve as the original 0.1 weight decay (the difference between v2.4 and v2.6 is only the weight decay, and since they yield the same curve v2.6 completely hides v2.4, rendering it invisible in the plot).
In conclusion, all of the runs (summarized in Figure 12) had the same outcome of eventual flattening of the validation loss and sometimes even increasing the loss. We did not observe that any particular change significantly improved the downstream evaluations and validation loss (Table 20).
v2 v2.1 v2.2 v2.3 v2.4 v2.5 v2 v2.1 v2.2 v2.3 v2.4 v2.5
step $_{\text{future}} val$$_{\text{past}}$
99300 8.43 8.60
115500 8.30 8.43
126600 8.34 8.24 8.31 8.40 8.32 8.56
131700 8.09 8.20 8.22 8.24
133200 8.35 8.08 8.38 8.22
137100 8.13 8.30
139200 8.14 8.32
143400 8.15 8.33
145800 8.16 8.32
[…]
99300 37.77 43.57
115500 38.79 43.10
126600 37.86 38.09 38.77 42.37 42.79 42.82
131700 38.76 38.51 43.02 43.49
133200 37.02 38.90 42.26 43.46
137100 38.71 44.02
139200 39.02 44.20
143400 38.98 43.20
145800 38.80 43.37
[…]
At this point, we had used 77% of our training data and were nearing the end of the budget we had set aside for training. Combined with all of these observations and initial promising results on the downstream benchmarks, we decided to end training despite not having gone through all of our training data. Another motivating factor was the possibility of using remaining unseen training data for subsequent runs of different styles of training and finetuning.
Based on this experience, we plan to explore different options in future experiments that have shown the potential to lead to more stable training for longer durations, including SwiGLU activations (Shazeer, 2020), RoPE embeddings (Su et al., 2021b), and normalization for queries and keys in the attention layers (Henry et al., 2020).